OpenStudy (anonymous):

who knows how to find the area of a circle? will medal first to answer whith the correct answer.

OpenStudy (anonymous):

A = π r 2

OpenStudy (anonymous):

The area of a circle is π (Pi) times the Radius squared, which is written: A = π × r2 Or, when you know the Diameter: A = (π/4) × D2 Or, when you know the Circumference: A = C2 / 4π

OpenStudy (anonymous):

a= 3.14* r squared

OpenStudy (anonymous):

(? = pi)

OpenStudy (anonymous):

thank you. :)

OpenStudy (anonymous):

np